OpenStudy (anonymous):

What is the equation of a circle with a center at (-4, -8) and a radius of 1.9?

OpenStudy (anonymous):

13.8

OpenStudy (anonymous):

That isn't a equation, but ok

OpenStudy (anonymous):

oh the equation of the circle is 3.9 in the iquisdent of 402 and 29 to 39

OpenStudy (rmrjr22):

(x - h)2 + (y - k)2 = r2

OpenStudy (rmrjr22):

(h,k) = center points

OpenStudy (anonymous):

Oh so I just plug the values in. thanks!

OpenStudy (rmrjr22):

Np :)

OpenStudy (rmrjr22):

u know u wanna click that best response button ;)

OpenStudy (anonymous):

Haha okayyyy I will

OpenStudy (rmrjr22):

Thanks, im officially A "life saver"

OpenStudy (anonymous):

What is the equation of a circle with a center at (-3, 6) and a radius of 7? Would this one be (x-3)^2 + (y-6)^2= 7^2?

jimthompson5910 (jim_thompson5910):

close, more like (x-(-3))^2 + (y-6)^2 = 7^2 that turns into (x+3)^2 + (y-6)^2 = 49

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (rmrjr22):

be careful with signs.. they get you alot

OpenStudy (anonymous):

SO if the center points are negative, does the minus sign in the equation become positive?

jimthompson5910 (jim_thompson5910):

yeah you have a sign flip

OpenStudy (rmrjr22):

yes... a negative * a negative = positive

OpenStudy (anonymous):

Got it! Do either of you know the equation formula for parabolas? What is the equation of a parabola with vertex (0, 0) and a focus at (2, 0)

OpenStudy (anonymous):

Or how to answer that question

OpenStudy (rmrjr22):

y = ax 2 + bx + c

OpenStudy (anonymous):

How would the values in the question go into that equation? For an example cause I have 10 problems like that :(

OpenStudy (rmrjr22):

y= a(x-h)2+k this is for the focus points, like the center of a circle. except its the center of the parabola

OpenStudy (anonymous):

Ah you def deserve the life saver badge. Thanks!